OpenStudy (anonymous):

what is the radius when the circumfrence is 22

myininaya (myininaya):

C=2*pi*r 22=2*pi*r 11=pi*r 11/pi=r

OpenStudy (saifoo.khan):

3.5units

OpenStudy (anonymous):

22 / 2pi

OpenStudy (anonymous):

radius = diameter/2 but the answer is 3.5